can the force on unit n pole betwn two isolated like poles of same strength be zero at any point??


I know the answer is the mid point betwen them, i am looking for the explanation to the answere...

Grade:Upto college level

1 Answers

Askiitians Expert Bharath-IITD
14 years ago
Dear Madhu, Let the two isolated poles of equal pole strengths are separated by distance L say. Now consider a point x assuming first pole is at origin. Then the net force experienced by a unit north pole placed at this point is given by the difference of the forces due to the two poles. Fnet= F1 - F2 = K{ (1/x)^2 - (1/L-x)^2 } Where K is a constant given by similar terms in both F1 and F2 i.e, {(munot/ 4* pi)*1* pole strength of the isolated pole}. If we equate it to Zero Then we would get x = L/2. Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best STUDENT NAME !!! Regards, Askiitians Experts Adapa Bharath